# Defense Against Prompt Injection Attack by Leveraging Attack Techniques

### v3

@@ 6 Conclusion
In this study, we explore the design of defense methods against prompt injection attacks by leveraging attack techniques, because of the similar design goals between the attack methods and the defense methods. Specifically, we design defense strategies based on the ignore attack, escape attack, fake completion attack, and fake completion attack with template. We evaluate our methods against both direct and indirect prompt injection attacks, comparing their performance to various training-free defense methods. The experimental results demonstrate that our defense methods outperform existing defense baselines, even decreasing the ASR to zero in some scenarios. What’s more, we observe that the stronger attack method can be utilized to build stronger defense method, paving the way for designing more effective defenses against more complex attacks in the future.
@@
In this paper, we propose defense methods inspired by existing attack strategies. However, since a benchmark of long queries for prompt injection research has not yet been established, we are unable to conduct a thorough investigation into how the truncation method addresses the long-query problem, as discussed in ablation study. As an alternative, we remove the original input instructions from existing benchmarks and provide approximate results. These results demonstrate the effectiveness of the proposed methods. Moreover, we do not employ gradient-based attack methods as defense methods, as previous studies have shown that their performance is not satisfactory. Finally, since our methods are based on prompt engineering, we focus on conducting comprehensive experiments to demonstrate their effectiveness, rather than providing a mathematical proof to explain why they work. This limitation can also be found in other prompt injection studies Liu et al. (2024b); Chen et al. (2024); Li et al. (2023); Hines et al. (2024), regardless of whether they are fine-tuning-based or prompt-engineering-based.
@@ A.2 Defense Baselines
A.3 Impact of deleting data content
A.4 Dataset Additional Information
AlpacaFarm is with license of Apache License 2.0 and dataset from Li et al. (2023) does not report the license. The datasets we use do not contain personal privacy information. AlpacaFarm is a simulation platform designed to facilitate research and development in learning from feedback, significantly reducing the typical costs. It aims to make research on instruction following and alignment more accessible Dubois et al. (2024). What’s more, dataset from Li et al. (2023) is constructed to evaluate the prompt inject attack. The datasets are open-sourced.
Defense Methods GPT-3.5-Turbo GPT-4o-Latest
None 32.69 50.48 32.69 88.46 87.50 65.86 92.78 63.46 100.00 100.00
Sandwich 13.94 17.30 8.65 4.32 42.30 15.86 29.32 10.09 5.76 37.98
Reminder 11.05 10.57 10.96 9.61 26.92 14.90 27.88 17.30 89.42 82.69
Isolation 22.59 39.42 24.51 43.26 77.40 52.40 83.17 52.88 94.71 99.51
Spotlight 16.34 31.73 13.46 15.38 71.15 19.71 45.67 15.38 47.11 68.75
Ours-Ignore 2.88 3.36 1.44 0.48 4.32 0.90 0.90 0.40 0.0 0.0
Ours-Fakecom 5.57 12.01 1.44 0.0 14.90 7.21 2.88 5.76 0.90 7.21

### v4

@@ 6 Conclusion
In this work, we explore the design of defense methods against prompt injection attacks by leveraging attack techniques, because of the similar design goals between the attack methods and the defense methods. We evaluate our methods against both direct and indirect prompt injection attacks, comparing their performance with various training-free defenses as well as fine-tuning-based approaches. The experimental results demonstrate that our defense methods outperform existing defense baselines, even decreasing the ASR to zero in some scenarios. What’s more, we observe that the stronger attack method can be utilized to build stronger defense method, paving the way for designing more effective defenses against more complex attacks in the future.
@@
In this paper, we propose defense methods inspired by existing attack strategies. However, since a benchmark of long queries for prompt injection research has not yet been established, we are unable to conduct a thorough investigation into how the truncation method addresses the long-query problem, as discussed in ablation study. As an alternative, we remove the original input instructions from existing benchmarks and provide approximate results. These results demonstrate the effectiveness of the proposed methods. Moreover, we do not employ gradient-based attack methods as defense methods, as previous studies have shown that their performance is not satisfactory. Finally, since our methods are based on prompt engineering, we focus on conducting comprehensive experiments to demonstrate their effectiveness, rather than providing a mathematical proof to explain why they work. This limitation can also be found in other prompt injection studies Liu et al. (2024b); Chen et al. (2024); Li et al. (2023b); Hines et al. (2024), regardless of whether they are fine-tuning-based or prompt-engineering-based.
@@
Acknowledgment
The work described in this paper was conducted in full or in part by Dr. Haoran Li, JC STEM Early Career Research Fellow, supported by The Hong Kong Jockey Club Charities Trust. We thank the authors of StruQ Chen et al. (2024) for providing the baseline code. We also sincerely appreciate Zhenran Xu, area chairs and reviewers for their valuable feedback and suggestions.
@@ A.2 Defense Baselines
A.3 Model Efficiency
Since LLMs are ultimately integrated into various applications, it is crucial to assess the overhead introduced by the defense method, particularly in terms of generating longer, but less informative, responses. To evaluate efficiency, we use the same QA dataset employed for assessing model utility. As demonstrated in Section 5.3.3, the utility results remain almost consistent across different defense strategies. Therefore, any additional overhead indicates an increase in uninformative content within the responses. Table 8 presents the efficiency of various defense methods. The average overhead of our defense method based on “Fakecom-t” and the defense based on “Escape attack” is slightly higher compared to the baseline with no defense, but it remains relatively low. The overhead for the other defense methods is nearly the same.
Defense Methods Llama3 Llama3.1 Qwen2
None 0.645 0.699 1.184
Sandwich 0.632 0.701 1.173
Instructional 0.630 0.701 1.169
Reminder 0.632 0.701 1.179
Isolation 0.631 0.702 1.171
Spotlight 0.644 0.734 1.193
Ours-Ignore 0.632 0.701 1.171
Ours-Escape 0.644 0.729 1.548
Ours-Fakecom 0.631 0.701 1.171
Ours-Fakecom-t 0.766 0.704 1.200
Table 8: The time cost across various defense methods. All the results are reported in seconds/item.